OpenStudy (anonymous):

Write an equation with a variable on both sides of the equal sign that has infinitely many solutions. Solve the equation and explain why it has an infinite number of solutions.

OpenStudy (anonymous):

I just dont know two equations that would give us the results that are infinite!

OpenStudy (anonymous):

thinking....

OpenStudy (anonymous):

help please i dont any of this tuff

OpenStudy (anonymous):

what math is it?

OpenStudy (anonymous):

eight grade math

OpenStudy (anonymous):

will this work? 2x + 6 = 4/2x + 12/6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah that actually a good example

OpenStudy (anonymous):

now all i need is the reason why

OpenStudy (anonymous):

please tell me why it is infitie

OpenStudy (anonymous):

do u know how to solve it? sorry having network problems

OpenStudy (anonymous):

no

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry kidd

OpenStudy (anonymous):

please tell me

OpenStudy (anonymous):

ok 2x + 6 = 4/2x + 12/2 Now you have to get variables on one side and numbers on the other Variables Numbers 2x + 6 = 4/2x + 12/2 -4/2x -4/2x 0x + 6 = 12/2 -6 -6 0 = 0 Since the answeer is the same on both sides of the equal sign, it has many solutions!

OpenStudy (anonymous):

Do u get it?

OpenStudy (anonymous):

yes now i do

OpenStudy (anonymous):

thank you nicole! :)

OpenStudy (anonymous):

bye guys! I am having trouble with my stupid laptop!!! I will try to log in later using the desktop!!!

OpenStudy (anonymous):

I needed help with that question too but i don't get it still?